# Efficient Online Inference of Vision Transformers by Training-Free Tokenization

### v1

@@ Abstract
The cost of deploying vision transformers increasingly represents a barrier to wider industrial adoption. Existing compression requires additional end-to-end fine-tuning or incurs a significant drawback to runtime, thus making them ill-suited for online inference. We introduce the Visual Word Tokenizer (VWT), a training-free method for reducing energy costs while retaining performance and runtime. The VWT groups patches (visual subwords) that are frequently used into visual words while infrequent ones remain intact. To do so, intra-image or inter-image statistics are leveraged to identify similar visual concepts for compression. Experimentally, we demonstrate a reduction in wattage of up to 19% with only a 20% increase in runtime at most. Comparative approaches of 8-bit quantization and token merging achieve a lower or similar energy efficiency but exact a higher toll on runtime (up to 2×2$\times$2 × or more). Our results indicate that VWTs are well-suited for efficient online inference with a marginal compromise on performance.
@@
Figure 1: Visualization of patch matching by our visual word tokenizer (${\mathcal{T}}_{inter}^{100}$ model). Patches that are matched with one another are indicated by identical colors. Higher patch matching is exhibited by the background rather than foreground object across the datasets. The patch matching of ${\mathcal{T}}_{inter}^{100}$ serves as a rudimentary form of image segmentation by grouping similar non-adjacent visual concepts.
Figure 2: Overview of the Visual Word Tokenizer. (a) An intra-image approach. In the forward pass, the pixel variance of the patches is computed with bottom-k values being masked. Grouped tokens are dropped after positional information is added. Note that ${\mathcal{V}}=1$ as only a single pixel variance feature is used. (b) An inter-image approach. A Bag-of-Visual Words is formed by clustering patches in the image space. In the forward pass, the minimum pairwise cosine distance between the patches and visual words is computed with values above the threshold being masked. Grouped tokens are averaged after positional information is added.
@@ 1 Introduction
In recent years, deep learning has seen continuous integration into a variety of systems worldwide. From coding to gaming, neural networks are increasingly deployed in online scenarios, where asynchronous requests are processed in real-time. However, due to the size and complexity of modern architectures, such models are costly to run in practice. Various methods have been proposed to improve model efficiency such as Knowledge Distillation [22], Pruning [21, 38], and Quantization [11]. However, many of these methods either require end-to-end fine-tuning to recover performance or significantly reduce runtime. In the field of Natural Language Processing (NLP), there is a growing trend towards improving efficiency via tokenization [18, 19, 9, 61, 40]. Newer large language models (LLMs) [53, 14] exhibit a noticeably larger vocabulary than their earlier counterparts [12, 42], thereby producing shorter sequences across various distributions. For computer vision, increasing interest is placed on reducing the cost of deploying the vision transformer (ViT) [13]. As image encoders in larger vision-language systems, ViTs are used to process images as fixed sets of tokens. Similar to downsampling in convolutional neural networks, most research [25, 30, 45, 37, 5, 4, 24] has focused on merging and/or pruning tokens in the intermediate layers to reduce computational overhead. Given the analogous architecture of the transformer across image and text modalities, our work looks instead at the idea of tokenization for efficiency by splitting an image into variable sets of tokens.
To introduce variability, we draw inspiration from subword tokenization algorithms used in NLP which follow the principle that common words should remain intact while infrequent words are broken down into meaningful subword units. Instead of a top-down approach – splitting words into subwords, our work for image data takes a bottom-up approach by grouping visual subwords (image patches) into visual words. We also twist the underlying principle: frequently used image patches should be grouped as they are more likely to describe common features or backgrounds while infrequent ones remain intact as they might carry task-relevant information. We propose two procedures to capture this principle. The first is an intra-image approach where patches with the lowest pixel variance within each image are grouped as they are more likely to represent uniform areas (e.g. backgrounds). The second is an inter-image approach, where basic features across multiple images such as colors or edges are discovered as visual words. Image patches are then grouped based on the similarity of these basic characteristics. Crucially, image patches that have distinct characteristics (i.e. high dissimilarity with any visual word) remain intact and form visual subwords. The code for our paper is publicly available [^1] https://github.com/wearepal/visual-word-tokenizer.
@@ 2 Related Work
Most works for improving the efficiency of ViTs have focused on reducing tokens in the intermediate layers by leveraging importance scores. In [30, 60, 25, 4], redundancy is addressed by fusing tokens. Both [45] and [50] opt to prune such tokens instead. Recent efforts [7, 8, 6, 24] attempt to combine the benefits of merging and pruning. In [54], an additional metric termed the energy score is used to better identify redundancy. Uniquely, [16] use inverse transform sampling to select important tokens. Most relevant to our work are [37] and [5]. The former assigns tokens to centroids via clustering, while the latter progressively merges tokens in a training-free manner.
@@
Our method also takes inspiration from works in NLP for efficient inference. Increasingly, the tokenizer’s vocabulary is specialized to reduce the token length of the given distribution. In [18], domain adaptation of the tokenizer by Fast Vocabulary Transfer (FVT) ensures fewer subword or character tokens are produced. [19] followed up by introducing n-grams for tokenization beyond the word-level boundary. Vocabulary specialization coupled with FVT has also been shown to accelerate code generation for modern LLMs [9]. Meanwhile, [61] analyzed the effectiveness of various vocabulary adaptation techniques for efficient cross-lingual inference. Recently, [40] leveraged hypernetworks for zero-shot tokenizer transfer of newly domain-adapted vocabularies.
Vector Quantization.
The idea of discretizing continuous distributions has been explored in many works, most recently for image generation. [62] leveraged clustering for learning a codebook that maps keypoint descriptors to discrete visual words. In [58, 3], discretization is applied as part of the modeling for ViTs. [55] learned discerete image representations by introducing the Vector Quantised-Variational Autoencoder (VQ-VAE) approach. [15] and [65] further improved upon the VQ-VAE by combining the expressiveness of transformers with an adversarial framework. Increasingly, vision-language models paired with codebooks conduct image synthesis autoregressively [44, 66, 35, 36, 52, 51, 49]. Lastly, [63] tackled disentangled representation learning and scene decomposition by tokenizing images into visual concepts.
3 Visual Word Tokenizer
In ViTs, tokenization is a process that splits an image into patches (tokens) which are then projected into a series of embeddings. The number of patches is typically fixed (e.g. $197$) based on the choice of architecture and hyperparameters. We seek to split the image into variable length inputs instead (i.e. certain images will use $80$ tokens, while others a $100$). This variability is induced by conventional text tokenizers [17, 48, 59, 26] for model efficiency. We propose to achieve this via visual words that group patches (visual subwords) based on some commonality in the image space. These visual words capture frequently used patches while retaining infrequent ones as is. A simple yet effective grouping can be done using either a criterion that looks at statistics of only one image (an intra-image approach) or across many images (an inter-image approach). Figure 2 summarizes the Visual Word Tokenizer (VWT).
3.1 An intra-image approach
The pixel variance of the image patches is the simplest criterion that can be used for grouping. In Figure 2, this approach only utilizes the deploy step by grouping the top-k patches with the lowest pixel variance while leaving the rest separate. To compress the grouped tokens, we opt to drop them as they tend to exhibit excessive homogeneity. We do not include the [CLS] token for dropping. This approach is inspired by [39] which aimed to reduce the training cost of OWLv2, an open-vocabulary object detector. Dropping patches with the lowest pixel variance removes padding areas and uniform backgrounds from the input mosaics of raw images, thus increasing training efficiency.
3.2 An inter-image approach
Inspired by text tokenizers and codebooks for image generation, we propose a variable length tokenizer that statistically discovers visual words across many images. The tokenizer consists of two steps: Pre-Process and Deploy
Pre-Process.
The Bag-of-Visual Words (BoVW) is a popular method for modeling images via discrete representations. In [62], k-means clustering is applied to keypoint descriptors from SIFT [34] to learn a fixed set of centroids. These centroids represent the vocabulary to which multiple descriptors are mapped in a process termed Vector Quantization (VQ). In our method, we adopt a variation of this framework by building the BoVW using patches from the image space. Our design choice is motivated by two main factors. First, we find keypoint descriptors to be costly for inference. In each forward pass, computing keypoints for each image would significantly increase runtime. Second, in our early experimentation, we observed that patches in the embedding space have little similarity to one another. Such issues were also described by [5], thus leading to their use of attention scores instead. Further justification for leveraging the image space is provided in Section B.1.
Given a dataset, we first patchify the images using the same patch size as the image encoder (e.g. 16 for ViT-B/16). We then cluster the patches via k-means to form the BoVW of a given vocabulary size, where each centroid represents a visual word. Patchification is done via basic tensor operations and not the pre-trained convolutional layer of the ViT to avoid projection into embeddings. We also execute this process in batches using the MiniBatchKMeans222from sklearn.cluster import MiniBatchKMeans algorithm due to memory constraints. Note that MiniBatchKMeans uses the Euclidean distance by design. Since clustering is done in the image space, the BoVW may be reused by other ViTs with the same patch size regardless of model type.
@@ 3.2 An inter-image approach
Dataset Model $Base$ $Q_{8}$ $ToME$ ${\mathcal{T}}_{intra}^{0.5}$ ${\mathcal{T}}_{inter}^{1000}$
Waterbirds CLIP 123.99 $\pm$ 3.07 73.89 $\pm$ 0.23 91.64 $\pm$ 0.23 110.93 $\pm$ 0.33 107.93 $\pm$ 0.21
CelebA 126.53 $\pm$ 0.75 74.47 $\pm$ 0.18 93.07 $\pm$ 0.13 114.37 $\pm$ 0.85 102.12 $\pm$ 0.20
MetaShift 123.30 $\pm$ 0.60 74.02 $\pm$ 0.43 90.96 $\pm$ 0.33 113.84 $\pm$ 2.40 105.78 $\pm$ 0.15
OpenImages (Com.) 135.05 74.80 94.63 114.62 112.67
OpenImages (Rare) 135.66 75.27 94.04 115.78 110.75
COCO BLIP 191.74 81.68 147.23 169.77 156.25
NoCaps 185.12 81.89 149.26 175.06 155.11
@@
Dataset Model $Base$ $Q_{8}$ $ToME$ ${\mathcal{T}}_{intra}^{0.5}$ ${\mathcal{T}}_{inter}^{1000}$
Waterbirds CLIP 8.48 $\pm$ 0.11 88.55 $\pm$ 0.10 17.04 $\pm$ 0.06 5.84 $\pm$ 0.03 9.73 $\pm$ 0.03
CelebA 8.33 $\pm$ 0.01 89.41 $\pm$ 3.30 17.25 $\pm$ 0.13 5.86 $\pm$ 0.05 9.54 $\pm$ 0.03
MetaShift 8.21 $\pm$ 0.09 89.48 $\pm$ 0.28 17.23 $\pm$ 0.34 6.01 $\pm$ 0.30 9.87 $\pm$ 0.05
OpenImages (Com.) 8.83 92.23 17.44 5.97 9.69
OpenImages (Rare) 8.60 86.00 17.04 6.19 10.02
COCO BLIP 12.45 57.11 15.06 8.10 9.10
NoCaps 12.44 57.61 15.08 8.04 9.28
@@
Table 2: Wattage and runtime per sample. Compared to $Base$, VWTs reduce wattages by up to 19% with ${\mathcal{T}}_{inter}^{1000}$. Although $Q_{8}$ and $ToME$ display lower wattages than VWTs, they induce significantly longer runtimes, especially the former. ${\mathcal{T}}_{intra}^{0.5}$ is consistently faster than $Base$, while ${\mathcal{T}}_{inter}^{1000}$ may increase runtimes by up to 20% at most (except on COCO and NoCaps).
Deploy.
Once the BoVW is formed, we now turn toward the process of sequence compression. One way of leveraging the BoVW would be to merge similar patches in the image space before projecting them into embeddings. However, such a naive approach will significantly degrade performance as doing so necessitates the initialization of a new linear layer for projection. To avoid this, we begin by patchifying and computing the pairwise cosine distance between the patches and BoVW. For each patch, we retain only the minimum distance. Unlike text, we are unable to obtain exact matches with images. As such, distances higher than a given threshold are masked out to ensure dissimilar patches are not merged. This allows us to ensure that infrequently used image patches remain intact. At this point, we have determined the groupings of similar patches via their connections to the same visual words. We then apply the pre-trained convolutional layer of the ViT on the original image to patchify and project it into a series of embeddings. Before merging, we ensure that positional information is added to the embeddings as we found it to work better than adding them later. Lastly, we average the embeddings element-wise based on the earlier defined groupings. We do not include the [CLS] token for merging.
4 Experiments
Consider a pre-trained image encoder $f\in{\mathcal{F}}$ with parameters $\theta\in\Theta$ that transforms inputs $x\in{\mathcal{X}}\subseteq\mathbb{R}^{d}$ to encodings $\hat{x}\in\hat{{\mathcal{X}}}\subseteq\mathbb{R}^{d}$. The encodings can then be mapped to labels $y\in{\mathcal{Y}}$ for some given task, be it classification or generation. More specifically, the ViT first transforms inputs $x$ to tokens $t_{1},\ldots,t_{N}$ before further processing by the attention layers. The number of tokens $N$ is a constant defined by $(\frac{I}{P})^{2}$, where $I$ and $P$ are the image and patch sizes, respectively. Let ${\mathcal{T}}$ be the VWT associated with a vocabulary $v\in{\mathcal{V}}$, where $v$ is a visual word learned from some dataset $D$. The tokenizer transforms the input $x$ into tokens $t_{1},\ldots,t_{M}$, where $M\ll N$. In our experiments, we seek to analyze the effect of ${\mathcal{T}}$ on online inference efficiency. We do so through the lens of (i) performance retention in visual recognition tasks of zero-shot classification, (ii) group robustness, and (iii) generative performance of visual captioning.
@@ 4 Experiments
$Base$ 79.06 $\pm$ 0.14 21.86 $\pm$ 0.32 89.61 $\pm$ 0.14 47.21 $\pm$ 2.41 95.31 $\pm$ 1.00 87.69 $\pm$ 3.08 70.48 63.36
$Q_{8}$ 79.94 $\pm$ 0.20 24.25 $\pm$ 0.09 89.73 $\pm$ 0.13 48.19 $\pm$ 3.24 95.23 $\pm$ 0.74 88.21 $\pm$ 3.87 70.48 63.19
$ToME$ 79.80 $\pm$ 0.15 27.05 $\pm$ 0.09 89.25 $\pm$ 0.15 28.52 $\pm$ 4.49 94.24 $\pm$ 0.43 87.37 $\pm$ 3.01 65.69 60.05
${\mathcal{T}}_{intra}^{0.5}$ 75.56 $\pm$ 0.06 31.00 $\pm$ 0.47 89.27 $\pm$ 0.12 50.93 $\pm$ 1.16 92.94 $\pm$ 0.35 86.15 $\pm$ 2.66 65.52 59.73
${\mathcal{T}}_{inter}^{1000}$ 79.19 $\pm$ 0.10 23.83 $\pm$ 0.27 90.79 $\pm$ 0.15 47.96 $\pm$ 4.98 93.94 $\pm$ 0.70 86.67 $\pm$ 2.35 66.15 60.56
(a) CLIP (image classification and group robustness)
@@
Table 3: Zero-shot image classification (CLIP), group robustness (CLIP) and captioning (BLIP). VWTs maintain average accuracy relative to $Base$ while improving robustness on Waterbirds. On COCO and NoCaps, VWTs retains a higher performance than $ToME$.
Dataset and Setting.
To study the effects of tokenization on (i) performance retention in visual recognition tasks of zero-shot classification, we conduct two sets of evaluations: 1) using one dataset for large-scale evaluation and 2) using three datasets for robustness analysis detailed below. We focus on a zero-shot setting by eschewing any end-to-end fine-tuning of $f$. For large-scale evaluation, we use the OpenImages v6 dataset [27]. Following [23], the test split is divided into common and rare subsets that consist of 57 224 and 21 991 images, respectively. The former has 214 classes, while the latter has 200. To perform classification, we compute the cosine similarity between an image embedding and the encoded text label [^3] The prefix ”a photo of a ” is also added to encode each text label..
To study performance retention and (ii) group robustness in image classification, we utilize three publicly available datasets: Waterbirds [57], CelebA [33], and MetaShift [29], that are used as benchmarks in robustness and fairness research [47, 32, 64]. To perform (zero-shot) classification, we compute the cosine similarity between an image embedding and the following encoded text labels3 for Waterbirds, CelebA, and MetaShift, respectively: {’landbird’, ’waterbird’}, {’non-blond’, ’blond’}, {’dog’, ’cat’}. The same data splits and spurious attributes as [32] and [64] are used for the former two and latter datasets, respectively. Specifically, the following spurious attributes are studied: habitation {’water’, ’land’} for Waterbirds, gender {’female’, ’male’} for CelebA, and context {’indoor’, ’outdoor’} for MetaShift. Further details on the subgroups are provided in Section A.1. We also compute the variance by shuffling samples among the data splits while maintaining subgroup proportions. We leverage attribute annotations not only for analyzing average and robust performance but also for the fairness of tokenization in Section B.4.
To analyze the effect of ${\mathcal{T}}$ on (iii) generative task of image captioning, we utilize the Karpathy test split of COCO dataset [31] and a validation set of NoCaps dataset  [1] following the setting in previous work [28].
Finally, to study inference efficiency, we utilize all datasets for the downstream visual tasks (i)-(iii) described above. We focus on the online setting where batch size is 1.

### v2

@@ Abstract
The cost of deploying vision transformers increasingly represents a barrier to wider industrial adoption. Existing compression techniques require additional end-to-end fine-tuning or incur a significant drawback to runtime, making them ill-suited for online (real-time) inference, where a prediction is made on any new input as it comes in. We introduce the Visual Word Tokenizer (VWT), a training-free method for reducing energy costs while retaining performance and runtime. The VWT groups visual subwords (image patches) that are frequently used into visual words while infrequent ones remain intact. To do so, intra-image or inter-image statistics are leveraged to identify similar visual concepts for sequence compression. Experimentally, we demonstrate a reduction in wattage of up to 25% with only a 20% increase in runtime at most. Comparative approaches of 8-bit quantization and token merging achieve a lower or similar energy efficiency but exact a higher toll on runtime (up to 100% or more). Our results indicate that VWTs are well-suited for efficient online inference with a marginal compromise on performance.
@@
Figure 1: Visualization of patch matching by our visual word tokenizer (${\mathcal{T}}_{inter}^{100}$ model). Patches that are matched with one another are indicated by identical colors. Higher patch matching is exhibited by the background rather than foreground across datasets. Patch matching serves as a rudimentary form of image segmentation by grouping similar non-adjacent visual concepts.
@@ 1 Introduction
In recent years, deep learning has seen continuous integration into a variety of systems worldwide. From coding to gaming, neural networks are increasingly deployed in online scenarios where asynchronous requests are processed in real-time. However, due to the size and complexity of modern architectures, such models are costly to run in practice. To address this, various methods have been proposed to improve model efficiency such as Knowledge Distillation [24], Pruning [22, 41], and Quantization [11]. However, many of these methods either require end-to-end fine-tuning to recover performance or significantly reduce runtime. In the field of Natural Language Processing (NLP), there is a growing trend towards improving efficiency via tokenization [18, 19, 9, 67, 43]. Newer large language models (LLMs) [58, 14] exhibit a noticeably larger vocabulary than their earlier counterparts [12, 45], thereby producing shorter sequences across various distributions. For computer vision, increasing interest is placed on reducing the cost of deploying the vision transformer (ViT) [13]. As image encoders in larger vision-language systems, ViTs are used to process images as fixed sets of tokens. Similar to downsampling in convolutional neural networks, most research [27, 32, 48, 40, 5, 4, 26] has focused on merging and/or pruning tokens in the intermediate layers to reduce computational overhead. Given the analogous architecture of the transformer across image and text modalities, our work looks instead at the idea of tokenization for efficiency by splitting an image into variable sets of tokens.
To introduce variability, we draw inspiration from subword tokenization algorithms [17, 51] used in NLP which follow the principle that common words should remain intact while infrequent ones are broken down into meaningful subword units. Instead of a top-down approach – splitting words into subwords, our work for image data takes a bottom-up approach by grouping visual subwords (image patches) into visual words. We also twist the underlying principle: frequently used patches should be grouped as they are more likely to describe common features while infrequent ones remain intact as they might carry task-relevant information. We propose two procedures to capture this principle. The first is an intra-image approach where patches with the lowest pixel variance within each image are grouped as they typically represent uniform areas (e.g. backgrounds). The second is an inter-image approach where basic features across multiple images such as colors or edges are discovered as visual words. Image patches are then grouped based on the similarity of these basic characteristics. Crucially, patches that have distinct characteristics (i.e. high dissimilarity with any visual word) remain intact and form separate visual subwords. The code for our paper is publicly available [^1] https://github.com/wearepal/visual-word-tokenizer.
@@ 2 Related Work
Most works for improving the efficiency of ViTs have focused on reducing tokens in the intermediate layers by leveraging importance scores. In Liang et al. [32], Xu et al. [66], Kong et al. [27], Bian et al. [4], redundancy is addressed by fusing tokens. Both Rao et al. [48] and Tang et al. [55] opt to prune such tokens instead. Recent efforts [7, 8, 6, 26] attempt to combine the benefits of merging and pruning. In Tran et al. [60], an additional metric termed the energy score is used to better identify redundancy. Uniquely, Fayyaz et al. [16] use inverse transform sampling to select important tokens. Most relevant to our work are Marin et al. [40] and Bolya et al. [5]. The former assigns tokens to centroids via clustering, while the latter progressively merges tokens layer-by-layer in a training-free manner [^2] Unlike Bolya et al. [5], we do not include Marin et al. [40] as one of our baselines due to a lack of code release..
Figure 2: Overview of the Visual Word Tokenizer. An intra-image approach. In the forward pass, the pixel variance of the patches is computed with the top-k lowest values being masked (deploy). The grouped tokens are dropped after positional information is added. Note that ${\mathcal{V}}=1$ as only a single pixel variance feature is used. An inter-image approach. A Bag-of-Visual Words is formed by clustering patches in the pixel space (pre-process). In the forward pass, the minimum pairwise cosine distance between the patches and visual words is computed with values above the threshold being masked (deploy). The grouped tokens are averaged after positional information is added.
@@
Our method also takes inspiration from efficient inference in NLP. Increasingly, the tokenizer’s vocabulary is specialized to reduce the input token length. In Gee et al. [18], domain adaptation of the tokenizer ensures fewer subword or character tokens are produced. Gee et al. [19] followed up by introducing n-grams for tokenization beyond the word-level boundary. In Dagan et al. [9], tokenizer specialization is also shown to accelerate the task of code generation with modern LLMs. Meanwhile, Yamaguchi et al. [67] analyzed the effectiveness of various vocabulary adaptation techniques for efficient cross-lingual inference. Recently, Minixhofer et al. [43] leveraged hypernetworks for zero-shot tokenizer transfer of newly domain-adapted vocabularies.
@@
Dataset Model $Base$ $Q_{8}$ $ToME$ ${\mathcal{T}}_{intra}^{0.5}$ ${\mathcal{T}}_{inter}^{100}$ ${\mathcal{T}}_{inter}^{1000}$ ${\mathcal{T}}_{inter}^{10000}$
Waterbirds CLIP 123.99 73.89 91.64 110.93 102.65 107.93 117.36
CelebA 126.53 74.47 93.07 114.37 96.31 102.12 115.46
MetaShift 123.30 74.02 90.96 113.84 98.98 105.78 114.70
OpenImages (Com.) 135.05 74.80 94.63 114.62 106.33 112.67 123.95
OpenImages (Rare) 135.66 75.27 94.04 115.78 104.73 110.75 123.05
COCO BLIP 191.74 81.68 147.23 169.77 144.71 156.25 171.84
NoCaps 185.12 81.89 149.26 175.06 141.20 155.11 170.38
@@
Dataset Model $Base$ $Q_{8}$ $ToME$ ${\mathcal{T}}_{intra}^{0.5}$ ${\mathcal{T}}_{inter}^{100}$ ${\mathcal{T}}_{inter}^{1000}$ ${\mathcal{T}}_{inter}^{10000}$
Waterbirds CLIP 8.48 88.55 17.04 5.84 9.88 9.73 9.91
CelebA 8.33 89.41 17.25 5.86 9.45 9.54 9.99
MetaShift 8.21 89.48 17.23 6.01 9.99 9.87 9.75
OpenImages (Com.) 8.83 92.23 17.44 5.97 9.64 9.69 10.23
OpenImages (Rare) 8.60 86.00 17.04 6.19 9.69 10.02 9.67
COCO BLIP 12.45 57.11 15.06 8.10 8.73 9.10 10.31
NoCaps 12.44 57.61 15.08 8.04 9.49 9.28 10.51
@@
Table 2: Wattage and runtime per sample. Compared to $Base$, VWTs reduce wattages by up to 25% with ${\mathcal{T}}_{inter}^{100}$. Although $Q_{8}$ and $ToME$ display lower wattages than VWTs, they induce significantly longer runtimes, especially the former. ${\mathcal{T}}_{intra}^{0.5}$ is consistently faster than $Base$, while ${\mathcal{T}}_{inter}^{10000}$ may increase runtimes by up to 20% at most (except on COCO and NoCaps).
Vector Quantization.
The idea of discretizing continuous distributions has been explored in many works, most recently for image generation. Yang et al. [68] leveraged clustering for learning a codebook that maps keypoint descriptors to discrete visual words. In Wu et al. [64] and Bao et al. [3], discretization is applied as part of the modelling for ViTs. Van Den Oord et al. [61] learned discrete image representations by introducing the Vector Quantised-Variational Autoencoder (VQ-VAE) approach. Esser et al. [15] and Yu et al. [71] further improved upon the VQ-VAE by combining the expressiveness of transformers with an adversarial framework. Increasingly, vision-language models paired with codebooks conduct image synthesis autoregressively [47, 72, 38, 39, 57, 56, 54]. Lastly, Yang et al. [69] tackled disentangled representation learning and scene decomposition by tokenizing images into separate visual concepts.
3 Visual Word Tokenizer
In ViTs, tokenization is a process that splits an image into patches (tokens) which are then projected into a series of embeddings. The number of patches is typically fixed (e.g. $197$) based on the choice of architecture. We seek to split an image into variable length inputs instead (i.e. certain images will use $80$ tokens, while others a $100$). This variability is induced by conventional text tokenizers [17, 51, 65, 28] for model efficiency. We propose to achieve this via visual words that group patches (visual subwords) based on some commonality in the pixel space. These visual words capture frequently used patches while retaining infrequent ones as is. A simple yet effective grouping can be done using either a criterion that looks at statistics of only one image (an intra-image approach) or across many images (an inter-image approach). Figure 2 summarizes the Visual Word Tokenizer (VWT).
@@ 3 Visual Word Tokenizer
$Base$ 79.06 21.86 89.61 47.21 95.31 87.69 70.48 63.36
$Q_{8}$ 79.94 24.25 89.73 48.19 95.23 88.21 70.48 63.19
$ToME$ 79.80 27.05 89.25 28.52 94.24 87.37 65.69 60.05
${\mathcal{T}}_{intra}^{0.5}$ 75.56 31.00 89.27 50.93 92.94 86.15 65.52 59.73
${\mathcal{T}}_{inter}^{100}$ 78.41 26.01 90.04 53.15 90.24 84.28 62.90 58.10
${\mathcal{T}}_{inter}^{1000}$ 79.19 23.83 90.79 47.96 93.94 86.67 66.15 60.56
${\mathcal{T}}_{inter}^{10000}$ 79.68 22.90 90.12 46.85 94.81 86.15 69.03 62.50
(a) CLIP (image classification and subgroup robustness)
@@
Table 3: Zero-shot (training-free) image classification (CLIP), subgroup robustness (CLIP) and captioning (BLIP). VWTs maintain average accuracy relative to $Base$ while improving robustness on Waterbirds. On COCO and NoCaps, VWTs retain a higher performance than $ToME$.
3.1 An intra-image approach
The pixel variance of the image patches is the simplest criterion that can be used for grouping. In Figure 2, this approach only utilizes the deploy step by grouping the top-k patches with the lowest pixel variance while leaving the rest intact. To compress the grouped tokens, we opt to drop them as they tend to exhibit excessive homogeneity. We do not include the [CLS] token for dropping. This approach is inspired by Minderer et al. [42] which aimed to reduce the training cost of OWLv2, an open-vocabulary object detector. Dropping patches with the lowest pixel variance removes padding areas and uniform backgrounds from the input mosaics of raw images, thereby increasing training efficiency.
3.2 An inter-image approach
Inspired by text tokenizers and codebooks, we propose a variable-length approach that statistically discovers visual words across many images. The tokenizer consists of two steps: Pre-Process and Deploy
Figure 3: Visualization of patch dropping by our visual word tokenizer (${\mathcal{T}}_{intra}^{0.5}$ model). Patches with the lowest pixel variance that are dropped are indicated in black. In most cases, the dropped patches correspond to the background object on average which is typically uninformative.
Pre-Process.
The Bag-of-Visual Words (BoVW) is a popular method for modeling images via discrete representations. In Yang et al. [68], k-means clustering is applied to keypoint descriptors from SIFT [37] to learn a fixed set of centroids. These centroids represent the vocabulary to which multiple descriptors are mapped in a process termed Vector Quantization (VQ). In our method, we adopt a variation of this framework by building the BoVW using patches within the pixel space. Our design choice is motivated by two main factors. First, we find keypoint descriptors to be costly for inference. In each forward pass, computing keypoints for each image would significantly increase runtime. Second, in our early experimentation, we observed that patches in the embedding space have little similarity to one another. Such issues were also described by Bolya et al. [5], thus leading to their use of attention scores instead. Further justification for leveraging the pixel space is provided in Section 4.6.
Given a dataset, we first patchify the images using the same patch size as the image encoder (e.g. 16 for ViT-B/16). We then cluster the patches via k-means to form the BoVW of a given vocabulary size, where each centroid represents a visual word. Patchification is done via basic tensor operations and not the pre-trained convolutional layer of the ViT to avoid projection into embeddings. We also execute this process in batches using the MiniBatchKMeans333from sklearn.cluster import MiniBatchKMeans algorithm due to memory constraints. Note that MiniBatchKMeans uses the Euclidean distance by design. Since clustering is done in the pixel space, the BoVW may be reused by other ViTs with the same patch size regardless of model type.
Deploy.
Once the BoVW is formed, we turn towards the process of sequence compression. One way of leveraging the BoVW would be to merge similar patches in the pixel space before projecting them into embeddings. However, such a naive approach will significantly degrade performance as the initialization of a new linear layer for projection is required. To avoid this, we begin by patchifying and computing the pairwise cosine distance between the patches and BoVW. For each patch, we retain only the minimum distance. Unlike text, we are unable to obtain exact matches with images. As such, distances higher than a given threshold are masked out to ensure dissimilar patches are not merged. At this point, we have determined the groupings of similar patches via their connections to the same visual words. We then apply the pre-trained convolutional layer of the ViT on the original image to patchify and project it into a series of embeddings. Before merging, we ensure that positional information is added to the embeddings as we found it to work better than adding them later. Lastly, we average the embeddings element-wise based on the earlier defined groupings. We do not include the [CLS] token for merging.
4 Experiments
Consider a pre-trained image encoder $f\in{\mathcal{F}}$ with parameters $\theta\in\Theta$ that transforms inputs $x\in{\mathcal{X}}\subseteq\mathbb{R}^{d}$ to encodings $\hat{x}\in\hat{{\mathcal{X}}}\subseteq\mathbb{R}^{\hat{d}}$. The encodings can then be mapped to labels $y\in{\mathcal{Y}}$ for some given task, be it classification or generation. More specifically, the ViT first transforms inputs $x$ to tokens $t_{1},\ldots,t_{N}$ before further processing by the attention layers. The number of tokens $N$ is a constant defined by $(\frac{I}{P})^{2}$, where $I$ and $P$ are the image and patch sizes, respectively. Let ${\mathcal{T}}$ be the VWT associated with a vocabulary $v\in{\mathcal{V}}$, where $v$ is a visual word learned from some dataset $D$. The tokenizer transforms the input $x$ into tokens $t_{1},\ldots,t_{M}$, where $M\ll N$. In our experiments, we seek to analyze the effect of ${\mathcal{T}}$ on online inference efficiency. We focus on the zero-shot setting by eschewing any form of end-to-end fine-tuning for $f$.